# ESG Rating Disagreement and Corporate Total Factor Productivity:Inference and Prediction

### v4

@@ 6 Conclusion and Implication
ESG rating disagreements present significant challenges to the improvement of corporate TFP, underscoring the urgent need for firms to address this issue. Mechanism analysis reveals that green innovation can exacerbate the negative effects of ESG rating disagreements on TFP. Additionally, these disagreements hinder TFP growth by intensifying financing constraints and weakening human capital. To mitigate these impacts, companies should focus on improving the quality of ESG information disclosure, carefully navigate the trade-offs between green innovation, financing needs, and strive to balance the interests of diverse stakeholders. Regulatory authorities, in turn, should strengthen the scrutiny of green innovation patents to ensure their quality and prevent superficial compliance efforts.
Heterogeneity analysis demonstrates that the negative effects of ESG rating disagreements on TFP are more pronounced in state-owned enterprises, non-capital-intensive firms, low-pollution industries, and high-tech enterprises. These entities should pay closer attention to ESG rating discrepancies, while regulators should provide targeted support to low-pollution and high-tech firms to foster sustainable development and technological innovation. Moreover, ESG rating disagreements serve as a critical predictor of TFP, with SHAP values highlighting their substantial and ongoing negative impact. This indicator can be leveraged by listed companies and regulatory agencies as a monitoring and early warning tool to identify potential risks to TFP and implement proactive measures.
Appendix A Appendix
@@ A.2 Correlation Coefficient
A.3 Robustness Tests and Endogeneity Treatment
@@
  • Note: TFP is t+1 period, just like the baseline regression setup.

### v5

@@ 6 Conclusion and Implication
ESG rating disagreement presents significant challenges to the improvement of corporate TFP, underscoring the urgent need for firms to address this issue. Mechanism analysis reveals that green innovation can exacerbate the negative effects of ESG rating disagreement on TFP. Additionally, these disagreement hinder TFP growth by intensifying financing constraints and weakening human capital. To mitigate these impacts, companies should focus on improving the quality of ESG information disclosure, carefully navigate the trade-offs between green innovation, financing needs, and strive to balance the interests of diverse stakeholders. Regulatory authorities, in turn, should strengthen the scrutiny of green innovation patents to ensure their quality and prevent superficial compliance efforts.
Heterogeneity analysis demonstrates that the negative effects of ESG rating disagreement on TFP are more pronounced in state-owned enterprises, non-capital-intensive firms, low-pollution industries, and high-tech enterprises. These entities should pay closer attention to ESG rating discrepancies, while regulators should provide targeted support to low-pollution and high-tech firms to foster sustainable development and technological innovation. Moreover, ESG rating disagreement serve as a critical predictor of TFP, with SHAP values highlighting their substantial and ongoing negative impact. This indicator can be leveraged by listed companies and regulatory agencies as a monitoring and early warning tool to identify potential risks to TFP and implement proactive measures.
We believe it would be worthwhile to further investigate how green innovation amplifies the negative impact of ESG rating divergence on TFP, as well as to integrate SHAP with microeconometric methods for more robust empirical analysis.
Appendix A Online Appendix
@@ A.2 Correlation Coefficient
  • Note: This is the matrix of linear correlation coefficients for all variables.
A.3 Robustness Tests and Endogeneity Treatments
@@
  • Note: This is the result of only front-loading $Dis$ compared to $TFP$ by one period. In this regression result, $Dis$ denotes ESG Rating Disagreement of $t-1$ year and $TFP$ denotes total factor productivity of $t$ year. Unless otherwise noted, $Dis$ and $TFP$ are consistent with the baseline regression settings.
@@
  • Note: This is the result of the regression with the addition of control variables, and the other settings are consistent with the baseline regression. $N.TFP_{province}$ represents the level of total factor productivity at the provincial level.
@@
  • Note: This is the result of an instrumental variable test where $TFP$ is the t+1 period, just like the baseline regression.
@@
  • Note: This is the result after PSM, matched on a 1:1 basis, with other settings not differing from the baseline regression.
@@ A.4 Optuna Parameter Range
  • Note: This table shows the range of hyperparameters Optuna searches for, automatically selecting the optimal hyperparameters for the final model configuration.